OpenStudy (anonymous):

Simplify the expression. 3.7 – 1.8 – 3.67 + 4.4 – 1.34

OpenStudy (anonymous):

3.7-1.8=1.9 1.9-3.67=-1.77 -1.77+4.4=2.63 2.63-1.34=1.29